# NESTful: A Benchmark for Evaluating LLMs on Nested Sequences of API Calls

### v2

@@ Abstract
The resurgence of autonomous agents built using large language models (LLMs) to solve complex real-world tasks has brought increased focus on LLMs’ fundamental ability of tool or function calling. At the core of these agents, an LLM must plan, execute, and respond using external tools, APIs, and custom functions. Research on tool calling has gathered momentum, but evaluation benchmarks and datasets representing the complexity of the tasks have lagged behind. In this work, we focus on one such complexity, nested sequencing, with the goal of extending existing benchmarks and evaluation. Specifically, we present NESTful, a benchmark to evaluate LLMs on nested sequences of API calls, i.e., sequences where the output of one API call is passed as input to a subsequent call. NESTful contains 1800+ nested sequences where all the function calls are executable. Experimental results on multiple models and settings show that the best-performing model on the dataset has a full sequence match accuracy of 25% and win-rate of 34% necessitating a large scope for improvement in the nested sequencing aspect of function calling. Our analysis of these results provides possible future research directions for the community, in addition to a benchmark to track progress. We have released the NESTful dataset under the Apache 2.0 license at https://github.com/IBM/NESTFUL.
@@ 1 Introduction
The significant reliance on LLMs’ function calling abilities led recent research to continuously improve this dimension of LLMs. On one hand, approaches to improve function calling have exploded Abdelaziz et al. (2024); Liu et al. (2024); Srinivasan et al. (2023); on the other hand, benchmarks and evaluations are lagging behind. For instance, BFCL v1 and v2 focused on evaluating single, multiple, and parallel function calling tasks for both non-executable and executable versions Patil et al. (2023). Works such as API-Blend Basu et al. (2024) complemented prior work by introducing granular task evaluation such as slot-filling, API-detection, and sequencing. BFCL v3 has progressed further into agentic use cases with multi-step and multi-turn function calling evaluation.
@@
In this paper, we present NESTful, a benchmark specifically designed to evaluate models on nested API calls and it contains over 1800 nested sequences. The data consists of: (1) user queries, (2) a catalog of APIs and their specifications, (3) the sequence of API calls and the corresponding parameters, and (4) the expected output response. The datasets are based on the MathQA Amini et al. (2019) and StarCoder2-Instruct Wei et al. (2024) datasets which are commonly used in the literature but are missing the executable component. Figure 1 shows the end-to-end evaluation pipeline of NESTful.
We evaluated the dataset on 15 state-of-the-art models from the literature and exposed the gaps of these models in handling complex function calling sequences. The best model did not exceed 25% full sequence match nor 33% on win rate metrics. Models struggled as the nesting got deeper and the data dependencies increased. To further advance research in this area, we will publicly release the NESTful dataset with executable Python implementation for each tool and the evaluation code for the models.
@@ 2 Related Work
Most of these existing works initially focused on basic function calling abilities that did not involve much complexity. Recent advances in enabling LLMs to handle complex multi-API interactions have introduced structured methods like Reverse Chain Zhang et al. (2024b). This approach employs backward reasoning to optimize multi-step API planning, allowing LLMs to effectively manage nested workflows by aligning intermediate steps with the final goal. Such methods highlight LLMs’ potential to perform efficient, target-driven planning in resource-constrained environments.
To evaluate and enhance these aforementioned approaches, numerous works released training and benchmarking data in service of API function calling, such as ToolLLM Qin et al. (2023); APIBench Patil et al. (2023); APIGen Liu et al. (2024); or API-BLEND (Basu et al., 2024). While those mostly focused on simpler tasks (e.g., single or multiple isolated API calling functionality), NesTools (Han et al., 2024) focused on more complex real-world applications. It was designed to evaluate the nested tool learning capabilities of LLMs, addressing the lack of nested sequencing related data in the existing benchmarks. Similarly, our work focuses on the subset of function calling tasks that need an interdependent sequence of API calls, a necessity in many real-world, multi-step problems.
Unlike NesTools which was created through a fully synthetic process, NESTful is derived from well-known existing datasets. Additionally, the average sequence length for NESTful is 4.36, compared to 3.04 for NesTools. Other function-calling benchmarks, such as SealTool Wu et al. (2024) and BFCL-v3222https://gorilla.cs.berkeley.edu/blogs/13_bfcl_v3_multi_turn.html, include some nested sequences but are smaller than NESTful as they are not specifically designed for nested function calls.
Finally, AgentBoard (Ma et al., 2024) is a comprehensive benchmark and open-source evaluation framework designed for assessing LLMs as general-purpose agents. It introduces fine-grained progress metrics and interactive visualization tools to analyze incremental advancements and multi-faceted agent behaviors, addressing the limitations of traditional success-rate-focused evaluations. This framework enhances interpretability and fosters the development of more robust LLM agents. In contrast, our NESTful benchmark is more focused on the evaluation of tool-augmented LLMs and calculates granular offline metrics (F1 score, accuracy, etc.) in addition to the win rate.
@@ 3 NESTful Dataset Curation
NESTful comprises of more than 1,800 instances designed for benchmarking tool-augmented LLMs, with a focus on nested sequencing. Each instance consists of (1) a user query, (2) a list of all available tools for the model to choose from, (3) the gold sequence of tools and their arguments needed to answer the user query, and (4) the final answer that should be obtained once the tools are executed. The dataset also contains corresponding Python code for every API in the library as well as a mechanism to run the input query via any LLM, execute the tools predicted by the LLM, and provide the final answer.
@@ 3.1 Nested Function Calling Data Schema
The NESTful data schema demonstrated in Figure 2 showcases the template used for representing the Input, Tool Library, Output, and Python Code. An important aspect of nested function calling is to enable a mechanism for tool reference; i.e. a subsequent tool call using that reference to access the output of the previous tool execution. To do so, we assign a unique variable name to each tool which distinctly identifies each tool, even when two identical tools with different arguments appear in the same sequence (parallel API calls). For example, in Figure  1, “rectangle_area” tool was assigned “label”: “\$var_1”. This allows the next tool “multiply” to use the output of “rectangle_area” as an argument: “arg_0”: “\$var_1.result\$”.
@@ 3.2.1 Mathematical Reasoning Data
For the first part of NESTful, we relied on datasets that test the model for nested function calling in the math domain. In particular, we build on MathQA Amini et al. (2019); a benchmark designed for evaluating the ability of AI models to solve mathematical word problems. It consists of questions that test numerical reasoning and problem-solving skills, requiring models to both understand the text of a word problem and perform mathematical operations to arrive at the correct solution. An example is shown in Figure 5.
@@
Figure 3: Sample specification for tools derived from MathQA
@@
Executable Code and Filtering
For each tool, we also generated its corresponding implementation in Python. This allows us to execute the nested call sequence and match the execution result with the gold answer. This ensures the correctness of the set of corresponding tools and the code execution too. We then filtered out any samples where we could not reproduce the gold answer from executing the nested tool calls. This process resulted in 1,415 test samples covering 40 tools with an average 5.1 tool calls per sample.
3.2.2 Coding Data
We also curated test examples based on generic Python functions from the StarCoder2-Instruct dataset Wei et al. (2024). This dataset has a total of 50K Python functions and covers a wide range of tools that can be used.
@@ 3.2.2 Coding Data
We leveraged StarCoder2-Instruct’s Python implementations and docstrings (such as in the example shown in Figure 4) to create API specifications. For each Python function, (1) we used an LLM (Mixtral 8x22B [^3] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1) to generate the possible Python types of the input and output arguments, and then (2) we validated and executed the function code itself to ensure it does not produce any errors. After both steps, we generated a corresponding JSON specification for each function documenting its input-output arguments. Figure 4 shows an example for “angle_rotation” function.
@@
To create input-output pairs, we synthetically generated 446 nested tool calling instances. To do so, we leveraged DiGiT [^4] https://github.com/foundation-model-stack/fms-dgt synthetic data generation framework where we manually curated 10 seed examples and used Mixtral-8x22B as the teacher model to generate the rest. We also built a validator to check the correctness of the produced samples w.r.t. ensuring the tools and parameters used are not hallucinated, required parameters are specified and there is at least one nested tool call in the output sequence.
@@ 4 Evaluation
Model Total Parameters One-shot ICL Three-shot ICL
F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate
xLAM-1b-fc-r 1B 0.35 0.16 0.13 0.04 0.03 0.49 0.20 0.15 0.06 0.04
xLAM-7b-fc-r 7B 0.54 0.32 0.18 0.10 0.13 0.49 0.35 0.23 0.18 0.19
MadeAgents/Hammer2.0-7b 7B 0.62 0.42 0.29 0.25 0.33 0.61 0.46 0.31 0.25 0.34
granite-3.0-8b-instruct 8B 0.50 0.27 0.15 0.05 0.07 0.52 0.35 0.22 0.15 0.16
llama-3-1-8b-instruct 8B 0.58 0.37 0.24 0.18 0.16 0.53 0.38 0.24 0.19 0.16
Team-ACE/ToolACE-8B 8B 0.51 0.23 0.13 0.00 0.00 0.51 0.27 0.14 0.00 0.00
llama-3-2-11b-instruct 11B 0.57 0.42 0.26 0.21 0.18 0.60 0.38 0.26 0.21 0.17
Granite-20B-FunctionCalling 20B 0.60 0.37 0.28 0.24 0.26 0.59 0.35 0.26 0.22 0.24
Mixtral-8x7B-Instruct-v0.1 46.7B 0.29 0.20 0.14 0.12 0.11 0.36 0.25 0.16 0.13 0.13
xLAM-8x7b-fc-r 46.7B 0.44 0.27 0.16 0.01 0.01 0.49 0.29 0.17 0.02 0.03
llama-3-1-70b-instruct 70B 0.35 0.18 0.10 0.00 0.01 0.33 0.16 0.10 0.00 0.03
llama-3-2-90b-instruct 90B 0.32 0.16 0.09 0.00 0.02 0.29 0.14 0.10 0.00 0.03
Mixtral-8x22B-Instruct-v0.1 141B 0.43 0.30 0.22 0.16 0.10 0.64 0.48 0.29 0.21 0.29
xLAM-8x22b-fc-r 141B 0.59 0.40 0.23 0.15 0.08 0.59 0.40 0.23 0.14 0.11
llama-3-1-405b-instruct-fp8 405B 0.21 0.20 0.13 0.10 0.08 0.26 0.25 0.15 0.11 0.14
Table 1: Evaluation results on NESTful on state-of-the-art LLMs with Direct Prompting technique. Models are sorted by their size. Experiments are done in one-shot and three-shot ICL settings. The best performance is highlighted in bold; the second best is underlined. Partial Sequence Accuracy (Part. Acc.) denotes the percentage of calling the correct API sequence (API names and arguments), whereas Full Sequence Accuracy (Full Acc.) counts the percentage of times where the model gets the entire sequence of APIs correctly. Both scores range from 0 to 1. We also report Win Rate which measures whether all the predicted APIs by the model are executable and lead to an exact match with the gold answer.
@@ 4.1 Baselines
We evaluated NESTful using 15 open-source models with sizes varying from 1B up to 405B parameters. This list includes the top models on Berkeley Function-Calling Leaderboard (BFCL) [^5] https://gorilla.cs.berkeley.edu/leaderboard.html. xLAM-1b-fc-r and xLAM-7b-fc-r (Liu et al., 2024) are fine-tuned function calling models based on DeepSeek-Coder-1.3B-instruct and DeepSeek-Coder-7B-instruct-v1.5. Similarly, xLAM model family also include xLAM-8x7b-fc-r [^6] https://huggingface.co/Salesforce/xLAM-8x7b-r and xLAM-8x22b-fc-r [^7] https://huggingface.co/Salesforce/xLAM-8x22b-r based on Mixtral 8x7B and Mixtral 8x22B, respectively. We also evaluate other competitive recent function calling models like Hammer2.0-7b [^8] https://huggingface.co/MadeAgents/Hammer2.0-7b and ToolACE-8B [^9] https://huggingface.co/Team-ACE/ToolACE-8B. We compare against various sizes of the latest LLAMA 3.1 models; particularly, llama-3-1-8b-instruct, llama-3-2-11b-instruct, llama-3-1-70b-instruct, llama-3-2-90b-instruct, llama-3-1-405b-instruct Dubey et al. (2024). We finally evaluate against other open models such as Granite-3.0-8b-instruct Granite Team (2024) and their tool calling variant Granite-20B-FunctionCalling Abdelaziz et al. (2024).

### v3

@@ Abstract
The resurgence of autonomous agents built using large language models (LLMs) to solve complex real-world tasks has brought increased focus on LLMs’ fundamental ability of tool or function calling. At the core of these agents, an LLM must plan, execute, and respond using external tools, APIs, and custom functions. Research on tool calling has gathered momentum, but evaluation benchmarks and datasets representing the complexity of the tasks have lagged behind. In this work, we focus on one such complexity, nested sequencing, with the goal of extending existing benchmarks and evaluation. Specifically, we present NESTful , a benchmark to evaluate LLMs on nested sequences of API calls, i.e., sequences where the output of one API call is passed as input to a subsequent call. NESTful contains 1800+ nested sequences where all the function calls are executable. Experimental results on a variety of models show that the best-performing model (GPT-4o) achieves a full sequence match accuracy of 28% and a win-rate of 60%, necessitating a large scope for improvement in the nested sequencing aspect of function calling. Our analysis of these results provides possible future research directions for the community, in addition to a benchmark to track progress. We have released the NESTful dataset under the Apache 2.0 license at https://github.com/IBM/NESTFUL.
@@ 1 Introduction
The significant reliance on LLMs’ function calling abilities led recent research to continuously improve this dimension of LLMs. On one hand, approaches to improve function calling have exploded Abdelaziz et al. (2024); Liu et al. (2024); Srinivasan et al. (2023); on the other hand, benchmarks and evaluations are lagging behind. For instance, BFCL v1 and v2 focused on evaluating single, multiple, and parallel function calling tasks for both non-executable and executable versions Yan et al. (2024). Works such as API-Blend Basu et al. (2024) complemented prior work by introducing granular task evaluation such as slot-filling, API-detection, and sequencing. BFCL v3 has progressed further into agentic use cases with multi-step and multi-turn function calling evaluation.
@@
In this paper, we present NESTful, a benchmark specifically designed to evaluate models on nested API calls with over 1800 nested sequences. It consists of: (1) user queries, (2) a catalog of APIs and their specifications, (3) the sequence of API calls and the corresponding parameters, and (4) the expected output response. The datasets are based on the MathQA Amini et al. (2019) and StarCoder2-Instruct Wei et al. (2024) datasets which are commonly used in the literature but are missing the executable component. Figure 1 shows the end-to-end evaluation pipeline of NESTful.
We evaluated the dataset on 19 state-of-the-art models from the literature and exposed the gaps of these models in handling complex function calling sequences. GPT-4o acheived the best performance, but did not exceed 28% full sequence match nor 60% on win rate metrics. Models struggled as the nesting got deeper and the data dependencies increased. To further advance research in this area, we will publicly release the NESTful dataset with executable Python implementation for each tool and the evaluation code for all models.
@@ 2 Related Work
Most of these existing works initially focused on basic function calling abilities that did not involve much complexity. Recent advances in enabling LLMs to handle complex multi-API interactions have introduced structured methods like Reverse Chain Zhang et al. (2024c). This approach employs backward reasoning to optimize multi-step API planning, allowing LLMs to effectively manage nested workflows by aligning intermediate steps with the final goal. Such methods highlight LLMs’ potential to perform efficient, target-driven planning in resource-constrained environments.
To evaluate and enhance these aforementioned approaches, numerous works released training and benchmarking data in service of API function calling, such as ToolLLM Qin et al. (2023); APIBench Patil et al. (2023); APIGen Liu et al. (2024); or API-BLEND (Basu et al., 2024). While these benchmarks focus on simpler or isolated API calls, NesTools (Han et al., 2024) and our work target more complex, real-world tasks involving interdependent, nested tool use. Unlike the fully synthetic NesTools, NESTful is built from established datasets and has longer average call sequences (4.36 vs. 3.04). Other benchmarks like SealTool Wu et al. (2024) and BFCL-v3 Yan et al. (2024) include some nesting but are smaller and not specifically designed for it.
ToolBench Xu et al. (2023) includes 205 nested samples from the Webshop and TableTop datasets, compared to over 1800 in NestFul. Moreover, ToolBench supports only 34 APIs, whereas NestFul features more than 900 unique functions. ShortcutsBench Shen et al. (2024) has multi-step tool calls but is tailored specifically to Apple Shortcuts feature, with evaluation data generated in a proprietary format that is difficult to interpret and does not follow a standard JSON structure. Finally, AgentBoard is a broad evaluation framework for general-purpose agents for multi-step planning tasks with fine-grained metrics and visual tools, while NESTful focuses specifically on tool-augmented LLMs with detailed offline metrics like F1 score and accuracy alongside win rate.
@@ 3 NESTful Dataset Curation
NESTful comprises of more than 1,800 instances designed for benchmarking tool calling in LLMs on nested sequencing. Each instance consists of (1) a user query, (2) a list of all available tools for the model to choose from, (3) the gold sequence of tools and their arguments needed to answer the user query, and (4) the final answer that should be obtained once the tools are executed. The dataset also contains corresponding Python code for every API in the library and a mechanism to run the input query via any LLM, execute the tools predicted by the LLM, and provide the final answer.
@@ 3.1 Nested Function Calling Data Schema
NESTful’s data schema, demonstrated in Figure 2, showcases the template used for representing the Input, Tool Library, Output, and Python Code. An important aspect of nested function calling is to enable a mechanism for tool reference; i.e. a subsequent tool call using that reference to access the output of the previous tool execution. To do so, we assign a unique variable name to each tool which distinctly identifies each tool, even when two identical tools with different arguments appear in the same sequence (parallel API calls). For example, in Figure  1, “rectangle_area” tool was assigned “label”: “\$var_1”. This allows the next tool “multiply” to use the output of “rectangle_area” as an argument: “arg_0”: “\$var_1.result\$”.
@@ 3.2 NESTful Data Domains
Figure 3: Sample specification for tools from MathQA
@@ 3.2.1 Mathematical Reasoning Data
For the first part of NESTful, we relied on datasets that test the model for nested function calling in the math domain. We build on MathQA Amini et al. (2019); a benchmark designed to evaluate AI models’ ability to solve mathematical word problems. It consists of questions that test numerical reasoning and problem-solving skills, requiring models to both understand the text of a word problem and perform mathematical operations to arrive at the correct solution.
@@
Executable Code and Filtering
For each tool, we also generated its corresponding implementation in Python. This allows us to execute the nested call sequence and match the execution result with the gold answer. It also ensures the correctness of the set of corresponding tools and the code execution too. We then filtered out any samples where we could not reproduce the gold answer from executing the nested tool calls. This process resulted in 1,415 test samples spanning 40 tools with an average 5.1 tool calls per sample.
3.2.2 Coding Data
We also curated test examples based on generic Python functions from the StarCoder2-Instruct dataset Wei et al. (2024). This dataset has a total of 50K Python functions and covers a wide range of tools that can be used. We started by collecting tool instructions and their Python implementations, followed by using Mixtral-8x22B to infer parameter type hints. Any functions that were syntactically incorrect or non-executable were filtered out. Next, a synthetic data generation pipeline was used to create instruction–nested call pairs using the valid seed tools and examples. This pipeline included a validator to ensure all parameters and tool names were accurate and complete. Finally, execution-based filtering was applied to verify that the generated samples produce the correct final output. We elaborate on each step in the following sections.
@@ 3.2.2 Coding Data
We leveraged StarCoder2-Instruct’s Python implementations and docstrings to create API specifications, see Figure 5 for an example. For each Python function, (1) we used Mixtral 8x22B [^2] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1 to generate the possible Python types of the input and output arguments, and (2) we validated and executed the function code itself to ensure it does not produce any errors. After both steps, we generated a corresponding JSON specification for each function documenting its input-output arguments. Figure 5 shows an example.
@@
To create input-output pairs, we leveraged DiGiT [^3] https://github.com/foundation-model-stack/fms-dgt synthetic data generation framework. DiGiT allows for defining various synthetic data generation pipelines given seed examples of input/output pairs. In particular, we used 10 seed examples and used Mixtral-8x22B as the teacher model. We also implemented a function calling validator that applies various heuristics to check the quality of the synthetic data, ensuring function calls adhere to the given specifications. In particular, we have validations to ensure the tools and parameters used are not hallucinated, required parameters are specified, and there is at least one nested tool call in the output sequence.
@@
3.3 Dataset Quality
Our benchmark builds on MathQA and StarCoder2-Instruct. MathQA is a well known mathematical reasoning dataset that was manually validated by humans, providing the input, correct sequence of math operations, and final answer. After converting it into a nested tool sequence, we further validate it by executing the sequence to ensure it produces the original gold answer. However, since the coding dataset is synthetically generated, we also implemented multiple automatic validation at various stages. This includes checking that nested tool sequences align with tool specifications and execute correctly to produce the expected output.
@@ 4 Evaluation
One-shot ICL Three-shots ICL
Model #Parameters F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate
xLAM-1b-fc-r 1B 0.19 0.08 0.09 0.00 0.01 0.22 0.09 0.09 0.03 0.02
xLAM-2-1b-fc-r 1B 0.41 0.13 0.13 0.00 0.00 0.43 0.12 0.13 0.00 0.00
xLAM-7b-fc-r 7B 0.49 0.17 0.15 0.00 0.03 0.55 0.23 0.23 0.15 0.14
xLAM-2-8b-fc-r 8B 0.48 0.15 0.14 0.00 0.01 0.47 0.17 0.15 0.04 0.04
Hammer2.0-7b 7B 0.56 0.24 0.21 0.07 0.16 0.61 0.30 0.29 0.22 0.25
Hammer2.1-7b 7B 0.10 0.05 0.05 0.01 0.01 0.16 0.10 0.11 0.08 0.08
Llama-3-1-8B-Instruct 8B 0.64 0.19 0.17 0.06 0.06 0.63 0.22 0.22 0.16 0.11
ToolACE-8B 8B 0.43 0.13 0.13 0.00 0.00 0.50 0.15 0.13 0.00 0.00
ToolACE-2-Llama-3.1-8B 8B 0.28 0.10 0.13 0.00 0.00 0.29 0.10 0.13 0.00 0.00
Granite-20B-FunctionCalling 20B 0.64 0.20 0.17 0.02 0.05 0.61 0.25 0.26 0.21 0.20
Mixtral-8x7B-Instruct-v0.1 46.7B 0.22 0.07 0.05 0.00 0.01 0.32 0.13 0.14 0.09 0.09
xLAM-8x7b-fc-r 46.7B 0.40 0.15 0.16 0.01 0.01 0.43 0.16 0.17 0.02 0.03
Llama-3-1-70B-Instruct 70B 0.41 0.19 0.15 0.04 0.09 0.33 0.17 0.15 0.07 0.11
Mixtral-8x22B-Instruct-v0.1 141B 0.49 0.21 0.17 0.06 0.07 0.65 0.29 0.28 0.21 0.23
xLAM-8x22b-fc-r 141B 0.53 0.21 0.22 0.12 0.03 0.50 0.23 0.25 0.17 0.06
Llama-3-1-405B-Instruct-fp8 405B 0.41 0.14 0.08 0.03 0.10 0.41 0.18 0.13 0.07 0.14
DeepSeek-V3 685B 0.69 0.36 0.27 0.09 0.43 0.69 0.42 0.37 0.29 0.60
GPT-4o (2024-08-06) UNK 0.73 0.41 0.38 0.28 0.59 0.74 0.41 0.38 0.28 0.60
Table 1: Evaluation results on NESTful on state-of-the-art LLMs with Direct Prompting technique. Models are sorted by their size. Experiments are done in one-shot and three-shot ICL settings. The best performance is highlighted in bold; the second best is underlined. Partial Sequence Accuracy (Part. Acc.) denotes the percentage of calling the correct API sequence (API names and arguments), whereas Full Sequence Accuracy (Full Acc.) counts the percentage of times where the model gets the entire sequence of APIs correctly. Both scores range from 0 to 1. We also report Win Rate, which measures whether all the predicted APIs by the model are executable and lead to an exact match with the gold answer.
@@ 4.1 Baselines
We extensively evaluated NESTful on 19 proprietary and open-source models, ranging in size from 1B to 685B parameters. This selection includes top tool-calling LLMs featured on the Berkeley Function-Calling Leaderboard (BFCL) Yan et al. (2024), as well as state-of-the-art models known for strong function-calling capabilities. Among the tool-calling models, we include the xLAM Zhang et al. (2024b); Liu et al. (2024), Hammer Lin et al. (2024), ToolAce Liu et al. model families, and Granite-20B-FunctionCalling Abdelaziz et al. (2024). We also evaluate a range of foundation models, including multiple sizes of LLAMA 3.1 Dubey et al. (2024), Mixtral [^4] https://huggingface.co/mistralai/, and DeepSeek-V3 Guo et al. (2024), and the state-of-the-art proprietary model GPT-4o Hurst et al. (2024). To explore how an agentic LLM performs on NESTful, we also include AgentLM-13B Zeng et al. (2023), which has been instruction-tuned using interaction trajectories from diverse agentic tasks.